Two airplanes depart simultaneously from an airport. One flies due North; the other flies due west at a rate 25m/h slower than the first airplane. After 3 hours, radar indicates that the airplanes are 1000miles apart. What is the ground speed of each airplane?
I got the equation: 18r^2-450r-994375=0
I need to solve the equation to get the answer, but I can't figure it out with the big numbers.
Thanks!!!

Answer by richwmiller(17219)

r*t=d
(3x)^2+(3*(x-25))^2=1000^2
Avoiding such large numbers
9*(2x^2-50x+625) = 1000000
divide by 9
The answer is not pretty.
x = 25/6*(3-sqrt(3191))
x = 25/6*(3+sqrt(3191))